OpenStudy (anonymous):

In the expression 3(3m+11) represents angle 1 and the expression 2(5m-10) -23 represents angle 2. What is the value of m? angle 1? and angle 2?

jimthompson5910 (jim_thompson5910):

what's the relationship between angles 1 and 2?

OpenStudy (anonymous):

they are adjacent

jimthompson5910 (jim_thompson5910):

are they complementary or supplementary?

OpenStudy (anonymous):

supplementary

jimthompson5910 (jim_thompson5910):

so angle 1 + angle 2 = 180

jimthompson5910 (jim_thompson5910):

which means that 3(3m+11) + 2(5m-10) -23 = 180

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Now solve for m to get 3(3m+11)+2(5m-10)-23=180 9m+33+10m-20-23=180 19m-10=180 19m=180+10 19m=190 m=190/19 m=10

jimthompson5910 (jim_thompson5910):

use this value of m to find the measures of the two angles angle 1: 3(3m+11) = 3(3*10+11) = 3(30+11) = 3(41) = 123 So angle 1 is 123 degrees You can plug m = 10 into the other expression, but we can exploit the fact that the two angles are supplementary So angle 2 = 180 - 123 = 57 degrees

OpenStudy (anonymous):

thanks